OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 6 cm2/min, find the rate at which the diameter decreases when the diameter is 8 cm.

OpenStudy (anonymous):

Assuming a spherical snowball with radius \(r\), its surface area \(A\) is given by \[A=2\pi r\]Differentiating with respect to time \(t\) yields \[\frac{dA}{dt}=2\pi \frac{dr}{dt}\]\(A\) decreases at a rate of \(6\text{ cm}^2/\text{min}\), which means \(\dfrac{dA}{dt}=-6\). You have enough info to solve for \(\dfrac{dr}{dt}\). The diameter is just twice the radius, i.e. \[d=2r\]so if you're looking for \(\dfrac{dd}{dt}\), it's just a matter of differentiating with respect to \(t\): \[\frac{dd}{dt}=2\frac{dr}{dt}\]